# Theme Article: NREL Special Issue

### v1

@@ Abstract
Adaptive Computing is an application-agnostic outer loop framework to strategically deploy simulations and experiments to guide decision making for scale-up analysis. Resources are allocated over successive batches, which makes the allocation adaptive to some objective such as optimization or model training. The framework enables the characterization and management of uncertainties associated with predictive models of complex systems when scale-up questions lead to significant model extrapolation. A key feature of this framework is the ability to explicitly utilize user-specified uncertainty priors, which we call model-specific local trust estimates, that are provided directly together with the problem specification and exploited in adaptive sampling strategies. A multi-fidelity model hierarchy is supported to allow trade-offs in accuracy and data acquisition cost while exploring the search space given a specified budget of potentially distributed, heterogeneous resources. We discuss application of this framework to problems in the renewable energy space, including biofuels production, material synthesis, perovskite crystal growth, and building electrical loads.
@@
Scale-up problems can be defined as the challenge of (1) translating a laboratory-scale experiment or small-scale simulation to a large-scale (e.g., industrial) application and (2) characterizing and managing the uncertainty introduced by scaling up the devices or processes involved thus leading to extrapolation beyond known data regimes. It is often inaccurate to directly extrapolate the results of laboratory-scale experiments to large-scale systems because new length and time scales can alter the balance of physical processes. For example, longer mixing times in a full-scale bioreactor might render it uneconomical at scale [1].
@@
Beyond differences in the modalities and representations of the underlying models, the drastic increase in the heterogeneity of computing resources has driven the development of a similar diversity in specialized workflows. These trends include central processing unit (CPU)-based high performance computing (HPC) resources being supplanted by heterogenous architectures including graphics processing units (GPUs), field-programmable gate arrays (FPGAs) and application-specific integrated circuits (ASICs), HPC centers being supplemented by flexible on-demand compute such as cloud-based services, and the proliferation of sensors and connected devices leading to increased opportunities for edge computing, i.e., discrete computing devices co-located with sensors to pre-process data and reduce data-load. Similarly there are many more real-time or autonomous user facilities, laboratories, and other data generating resources across the Department of Energy (DOE) [2], which bring their own challenges with regard to data management, resource management, cross comparison of results, and making decisions within a limited time frame (especially when steering or guiding experiment with computation).
In this paper we detail a computational outer-loop framework called Adaptive Computing (AC) to support decision making for scale-up analysis of renewable energy technologies, where data acquisition is constrained to a specified budget. This framework combines multi-fidelity surrogate modeling, optimal resource management across a diverse set of resources, and a novel scheme for incorporating and exploiting model-specific local trust estimates together with uncertainty estimation for adaptive sampling decisions. Users are able to specify uncertainty priors that vary across the sample space representing the uncertainty due to modeling assumptions and to restrict types of simulations or experiments to be used in feasible or trusted parameter regimes using constrains in the data acquisition optimization. These two capabilities make AC uniquely capable for addressing problems around technology scale-up, where data extrapolation plays a central role. Additionally, the modularity of the framework will enable communication across traditionally disciplinary silos (e.g., experimentalists, computational modelers, and HPC specialists).
@@
For a given problem the optimal surrogate model choice depends on the absolute cost of data acquisition and expected smoothness of the response. While GPs are robust with limited training data, their expressiveness is limited for complex or non-smooth mappings; in contrast NNs are much more data-hungry but can be universal function approximators.
@@
Uncertainty management begins with accurate quantification and propagation of epistemic and aleatoric uncertainties. Epistemic uncertainty can arise both from the individual models in a fidelity hierarchy (due to undersampling) and from the bridging functions that connect them (e.g., due to model bias). When using GPs as the underlying surrogate model, epistemic uncertainty is automatically quantified through samples in the latent Gaussian space; similarly other surrogate model classes can be used to manage epistemic uncertainty directly through, e.g., ensemble-based approaches. Aleatoric uncertainty can arise from inherent stochasticity in a simulation or experimental irrepeatability due to imprecise measurements or uncontrolled disturbances.
@@
Our approach is to leverage domain expertise to provide additional information about the trust one has about the validity of different models for specific regimes. For example, the high-fidelity model in the hierarchy for a fluid system could be based on a continuum representation operating well for low Knudsen numbers $Kn$, the non-dimensional number which tends toward zero in the continuum limit. Errors for the flow rate in internal flows are expected to scale with $c_{1}+c_{2}Kn+c_{3}Kn^{2}$ [5], where $c_{\cdot}$ are model constants. The user can provide this known deficiency of the Navier-Stokes for weakly rarefied flows by specifying this error scaling in the uncertainty prior for all continuum models in the multi-fidelity modeling hierarchy.
We call this user-specified prior a model-specific local trust estimate because it leverages physical insight (domain expertise) about the trustworthiness of a specific type of modeling assumption and it is a function of where the model is queried in the sample space, e.g., at which value of $Kn$. Similar trust estimates can be made for some experimental measurement techniques. For numerical methods, trust can depend on spatial and temporal resolutions (i.e., model hyperparameters); it can be smoothly varying, or discontinuous due, for example, to hidden constraints that prevent data acquisition in certain regions of parameter space. This trust model can then be exploited, not only to guide data acquisition, but to provide essential information for decision making given the outputs obtained from AC.
@@
Given the multi-fidelity surrogate with quantified uncertainty, we define a sampling strategy to identify new candidate function evaluations in the sample space to achieve the outer-loop goal. The first ingredient is selecting a criterion for the benefit $r_{m\ell}$ of computing a simulation, which is referred to as an acquisition function. For example, for optimization problems, the expected improvement (EI) acquisition function is commonly used with GPs [6]. This function balances local and global search by combining predicted function values with their corresponding uncertainty and trust estimates provided by the GP. Other acquisition functions exist such as probability of improvement, optimizing a confidence bound, using weighted sums of predicted function values and distance to already evaluated points, etc. Each has advantages and disadvantages [7].
In many multi-fidelity applications, we not only choose where to sample but also which fidelity level to evaluate while also optimizing over limited available resources. Moreover, sampling decisions may also be constrained by feasibility of an experiment or simulation at a given point in parameter space. AC’s sampling strategy thus exploits predictions from the multi-fidelity surrogate, uncertainty and trust estimates, and fidelity-specific feasibility constraints to decide the most promising points in the parameter space, and an optimization over the available resources allows optimal distribution of the evaluations. The goal is to arrive at the best possible solutions within the limited budget and to provide uncertainty estimates together with the solution to enable decision making.
An essential piece of the sampling strategy is batching, which permits parallel function evaluations. Here, the sampling strategy determines the next candidate for evaluation. Rather than immediately conducting this evaluation, it is added to a queue of points to be evaluated. The expected value from the surrogate model is treated as a placeholder value for the corresponding result, and the surrogate is retrained using this value. This algorithm is applied successively to develop a batch of $M_{\ell}$ candidate input locations for each fidelity level $\ell$. There is a trade-off between the benefits of frequently updating the inputs to the sampling strategy and the wall-clock efficiency of parallel sample evaluation that must be factored into the ideal batch size for a given application.
@@
The solution to the previous optimization problem is a set of cases (i.e., tasks) to run on computational or experimental hardware. (See the “data acquisition" step of Figure 1.) The orchestration of these tasks and their data is handled using the Hybrid Environment Resources and Operations (HERO) framework [8], which is a scalable, low-latency, and distributed task queue, capable of interfacing with the large number of heterogenous resources that may be necessary for a given problem (e.g., HPC resources with or without GPUs, a distributed cloud environment, edge devices such as network-attached microcontrollers, or experiment controllers). This scheme allows asynchronous execution across resources with vastly different time-scales and latencies, while simplifying the execution of on-demand model re-training based on non-blocking data availability.
@@
An engineering design problem can often be framed as an optimization over allowable ranges of design parameters such as material compositions and geometrical specifications. In the context of scale-up, design optimization has additional uncertainties since design decision must be made from a limited number of simulations of large-scale systems and many small-scale simulations and experiments. For illustration, we consider design optimization for an economically viable design of a chemical reactor for processing second-generation biofuels, which can be generated from waste (lignin- and cellulose-rich) biomass such as the non-edible parts of crops.
@@
The AC outer loop for this application is the design optimization problem. AC selects the input values (design specifications) to be simulated next based on surrogate-based optimization of some design objective. Specifically, enzymatic hydrolysis (EH) simulations are conducted using an National Renewable Energy Laboratory (NREL) software package called Vebio, a partial differential equation (PDE)-based computational fluid dynamics (CFD) code [9]. The default EH setup is used with the exception of two input design parameters, the solid mass fraction of xylan and the enzymatic load, which are optimized by AC to maximize the glucan concentration.
@@
Autonomous Laboratory Processes Design: Guided Synthesis of gallium nitride (GaN)
@@
The synthesis of a GaN thin film using metal modulated epitaxy (MME) encapsulates this difficulty. Though films have been grown with this method, controlling the vast arrays of associated process parameters (e.g. shutter timing, flux ratios, temperatures) and their influence on Ga droplet formation and film quality makes optimizing this process experimentally very expensive and difficult. Furthermore, while simulations can model this process and reproduce overall trends in, for example, defect discontinuity with shutter timing, there are substantial biases in the specific values in comparison with experiment [10]. This prohibits direct optimization of experimental parameters purely using simulation.
@@
We cast this problem into the AC framework by considering molecular dynamics (MD) simulations and experimental results as separate fidelities, with an outer loop problem of optimizing the process parameters for material quality. We then use the AC framework to learn the underlying connection between the simulation and experiment, while taking into account experimental uncertainty with limited samples.
@@
For illustration, at NREL, we are using the Adaptive Mesh and Algorithm Refinement (AMAR) approach. It allows for active coupling between mesh-based PDE solvers at the continuum level along with particle-based solvers at the molecular level. The computational domain capturing the physics at the reactor and device levels is discretized using a background mesh. At locations of governing physical phenomena, e.g., catalyst/electrode surfaces, the mesh is refined using adaptive mesh refinement (AMR) and the continuum solver is replaced with a different physical approach (e.g., kinetic Monte-Carlo (KMC) simulations), referred to as algorithm refinement. Information is transferred back and forth between the continuum and particle solvers in real-time. This has been used to model the dynamics inside a chemical vapor deposition reactor for perovskite crystal growth. The continuum flow solver transports species to a substrate. Several parallel KMC simulations are then initiated for each location on the substrate with its corresponding gas phase compositions, pressure and temperature. The net flux of species at the surface is coupled to the continuum as a boundary condition.
@@
The AC framework can also be used for engineering optimization for the design of the macro-scale conditions in the chemical vapor deposition reactor (e.g., ambient pressure, temperature, species composition, jet configuration), which determine the quality of the perovskite crystal grown at the substrate (e.g., number of defects, density, layer compositions, roughness). The AC framework is also useful for managing the uncertainty arising from coupling a deterministic continuum model with a stochastic particle model. Tracking the effect of the resulting uncertainties on the quantity of interests (QOIs) (e.g., crystal layer growth, response of the flow at the substrate) is of vital importance.
@@ Acknowledgements
We would like to acknowledge the domain experts who supported the integration of AC with our application areas. Nick Wunder and Monte Lunacek developed the HERO framework and aided with the integration with AC. Ethan Young and Hari Sitaraman developed the Vebio code and supported with our application in biofuel synthesis. Ross Larsen and Hari Sitaraman performed the AMAR simulations coupling KMC to the continuum flow. Davi Febba Marcelo, Stephen Schaefer, and Andriy Zakutayev developed the experimental infrastructure for the material synthesis portion, while Aidan Raver and Garritt Tucker provided simulation support.
@@
Kevin Patrick Griffin is Postdoctoral Research at NREL in Golden, CO, 80401, USA. His research interests include data-driven modeling, turbulence, and CFD. Dr. Griffin received his Ph.D.  in Mechanical Engineering from Stanford University. Contact him at kevin.griffin@nrel.gov.
@@
Jibonananda Sanyal is a Group Manager for Hybrid Energy Systems at [NREL]. His research interests include renewable energy, end-uses of energy, and applications of computing. Sanyal received his PhD in Computer Science from Mississippi State University. He is a Senior Member of IEEE. Contact him at jibo.sanyal@nrel.gov.

### v2

@@ Abstract
Adaptive Computing is an application-agnostic outer loop framework to strategically deploy simulations and experiments to guide decision making for scale-up analysis. Resources are allocated over successive batches, which makes the allocation adaptive to some objective such as optimization or model training. The framework enables the characterization and management of uncertainties associated with predictive models of complex systems when scale-up questions lead to significant model extrapolation. A key advancement of this framework is its integration of multi-fidelity surrogate modeling, uncertainty management, and automated orchestration of various computing and experimentation resources into a single integrated software package. This enables efficient multi-fidelity modeling across multiple computing resources by incorporating real-world constraints such as relative queue times and throughput on individual machines into the multi-fidelity sampling decision. We discuss applications of this framework to problems in the renewable energy space, including biofuels production, material synthesis, perovskite crystal growth, and building electrical loads.
@@
Scale-up problems can be defined as the challenge of (1) translating a laboratory-scale experiment or small-scale simulation to a large-scale (e.g., industrial) application and (2) characterizing and managing the uncertainty introduced by scaling up the devices or processes involved thus leading to extrapolation beyond known data regimes. It is often inaccurate to directly extrapolate the results of laboratory-scale experiments to large-scale systems because new length and time scales can alter the balance of physical processes. For example, longer mixing times in a full-scale bioreactor might render it uneconomical at scale.
@@
Beyond differences in the modalities and representations of the underlying models, the drastic increase in the heterogeneity of computing resources has driven the development of a similar diversity in specialized workflows. These trends include central processing unit (CPU)-based high performance computing (HPC) resources being supplanted by heterogenous architectures including graphics processing units (GPUs), field-programmable gate arrays (FPGAs) and application-specific integrated circuits (ASICs), HPC centers being supplemented by flexible on-demand compute such as cloud-based services, and the proliferation of sensors and connected devices leading to increased opportunities for edge computing, i.e., discrete computing devices co-located with sensors to pre-process data and reduce data-load. This diversity in computing architectures and workflows often lends itself to application-specific software for surrogate modeling and uncertainty management, limiting software reusability. Similarly, there are many real-time or autonomous user facilities, laboratories, and other data generating resources across the Department of Energy (DOE), which bring their own challenges with regard to data management, resource management, cross comparison of results, and making decisions within a limited time frame (especially when steering or guiding experiment with computation).
In this paper we detail a computational outer-loop framework called Adaptive Computing (AC) to support decision making for scale-up analysis, where data acquisition is constrained to a specified budget. AC integrates multi-fidelity surrogate modeling, adaptive sampling, and automated resource management to coordinate computational and experimental resources for applications in optimization, multiscale modeling, and control. The result is a flexible and modular tool that accommodates real-world HPC constraints, including resource schedulers, queue wait times, and heterogeneous computing environments. This integration enables more effective decision-making for large-scale simulations, ensuring that computational resources are allocated adaptively based on both model fidelity and system and resources availability.
@@
For a given problem the optimal surrogate model choice depends on the absolute cost of data acquisition and expected smoothness of the response. While GPs are robust with limited training data, their expressiveness is limited for complex or non-smooth mappings; in contrast NNs are much more data-hungry but can be universal function approximators. There are also projection-based (intrusive/non-intrusive) model reduction methods such as the proper orthogonal decomposition. See [1] for several recent developments. Model reduction methods can be efficient for representing high-dimensional data, but these methods are typically best suited for data that is structured in the input space. Also, these methods typically do not inherently provide uncertainty estimates. In this work, we focus on more general input-output black-box machine-learning techniques that are sufficiently robust to support lower dimensional quantities of interest with sparse data and that provide a mechanism for uncertainty quantification.
@@
Uncertainty management begins with accurate quantification and propagation of epistemic and aleatoric uncertainties [4]. Epistemic uncertainty can arise both from the individual models in a fidelity hierarchy (due to undersampling) and from the bridging functions that connect them (e.g., due to model bias). When using GPs as the underlying surrogate model, epistemic uncertainty is automatically quantified through samples in the latent Gaussian space; similarly other surrogate model classes can be used to manage epistemic uncertainty directly through, e.g., ensemble-based approaches. Aleatoric uncertainty can arise from inherent stochasticity in a simulation or experimental irrepeatability due to imprecise measurements or uncontrolled disturbances.
@@
It is often useful to leverage domain expertise to provide additional information about the trust one has about the validity of different models for specific regimes. For example, the high-fidelity model in the hierarchy for a fluid system could be based on a continuum representation operating well for low Knudsen numbers $Kn$, the non-dimensional number which tends toward zero in the continuum limit. Errors for the flow rate in internal flows are expected to scale with $c_{1}+c_{2}Kn+c_{3}Kn^{2}$ [5], where $c_{\cdot}$ are model constants. The user can provide this known deficiency of the Navier-Stokes for weakly rarefied flows by specifying this error scaling in the uncertainty prior for all continuum models in the multi-fidelity modeling hierarchy.
This model-specific local trust estimate is called a domain-expert-informed prior [4] because it leverages physical insight (domain expertise) about the trustworthiness of a specific type of modeling assumption and it is a function of where the model is queried in the sample space, e.g., at which value of $Kn$. Similar trust estimates can be made for some experimental measurement techniques. For numerical methods, trust can depend on spatial and temporal resolutions (i.e., model hyperparameters); it can be smoothly varying, or discontinuous due, for example, to hidden constraints that prevent data acquisition in certain regions of parameter space. This trust model can then be exploited, not only to guide data acquisition, but to provide essential information for decision making given the outputs obtained from AC.
@@
Given the multi-fidelity surrogate with quantified uncertainty, we define a sampling strategy to identify new candidate function evaluations in the sample space to achieve the outer-loop goal. The first ingredient is selecting a criterion for the benefit $r_{m\ell}$ of computing a simulation, which is the acquisition function value divided by the cost of the simulation. For example, for optimization problems, the expected improvement (EI) acquisition function is commonly used with GPs [6]. This function balances local and global search by combining predicted function values with their corresponding uncertainty and trust estimates provided by the GP. Other acquisition functions exist such as probability of improvement, optimizing a confidence bound, using weighted sums of predicted function values and distance to already evaluated points, etc. Each has advantages and disadvantages [7].
In many multi-fidelity applications, we not only choose where to sample but also which fidelity level to evaluate while also optimizing over limited available resources. Moreover, sampling decisions may also be constrained by feasibility of an experiment or simulation at a given point in parameter space. AC’s sampling strategy thus exploits predictions from the multi-fidelity surrogate, uncertainty and trust estimates, and fidelity-specific feasibility constraints to decide the most promising points in the parameter space, and an optimization over the available resources allows optimal distribution of the evaluations. The goal is to arrive at the best possible solutions within the limited budget and to provide uncertainty estimates together with the solution to enable decision making. The resource budget and wall-clock limit per iteration, $B_{i}$ and $T_{i}$, respectively, can be set as constants or estimated dynamically using look-ahead strategies to choose the batch size using estimates of future needs and the remaining budget [8]
An essential piece of the sampling strategy is batching, which permits parallel function evaluations [9]. Here, the sampling strategy determines the next candidate for evaluation. Rather than immediately conducting this evaluation, it is added to a queue of points to be evaluated. The expected value from the surrogate model is treated as a placeholder value for the corresponding result, and the surrogate is retrained using this value. This algorithm is applied successively to develop a batch of $M_{\ell}$ candidate input locations for each fidelity level $\ell$. There is a trade-off between the benefits of frequently updating the inputs to the sampling strategy and the wall-clock efficiency of parallel sample evaluation that must be factored into the ideal batch size for a given application.
@@
The solution to the previous optimization problem is a set of cases (i.e., tasks) to run on computational or experimental hardware. (See the “data acquisition" step of Figure 1.) The orchestration of these tasks and their data is handled using the Hybrid Environment Resources and Operations (HERO) framework, which is a scalable, low-latency, and distributed task queue, capable of interfacing with the large number of heterogenous resources that may be necessary for a given problem (e.g., HPC resources with or without GPUs, a distributed cloud environment, edge devices such as network-attached microcontrollers, or experiment controllers). This scheme allows asynchronous execution across resources with vastly different time-scales and latencies, while simplifying the execution of on-demand model re-training based on non-blocking data availability.
@@
An engineering design problem can often be framed as an optimization over allowable ranges of design parameters such as material compositions and geometrical specifications. In the context of scale-up, design optimization has additional uncertainties since design decision must be made from a limited number of simulations of large-scale systems and many small-scale simulations and experiments. For illustration, we consider design optimization for an economically viable design of a chemical reactor for processing second-generation biofuels, which can be generated from waste (lignocellulosic) biomass such as the non-edible parts of crops.
@@
The AC outer loop for this application is the design optimization problem. AC selects the input values (design specifications) to be simulated next based on surrogate-based optimization of some design objective. Specifically, enzymatic hydrolysis (EH) simulations are conducted using an National Renewable Energy Laboratory (NREL) software package called Vebio, a Virtual engineering framework with several zero and multidimensional models for biomass conversion processes [10]. EH is a process by which polymeric cellulose and hemicellulose molecules in biomass are broken down into monomeric sugars. The default EH setup is used with the exception of two input design parameters, the solid mass fraction of xylan (hemicellulose) and the enzyme concentration (g/L), which are optimized by AC to maximize the glucan (sugars) concentration.
@@
Autonomous Laboratory Processes Design: Guided Synthesis of GAN
@@
The synthesis of a gallium nitride (GaN) thin film using metal modulated epitaxy (MME) encapsulates this difficulty. Though films have been grown with this method, controlling the vast arrays of associated process parameters (e.g. shutter timing, flux ratios, temperatures) and their influence on Ga droplet formation and film quality makes optimizing this process experimentally very expensive and difficult. Furthermore, while simulations can model this process and reproduce overall trends in, for example, defect discontinuity with shutter timing, there are substantial biases in the specific values in comparison with experiment [11]. This prohibits direct optimization of experimental parameters purely using simulation.
@@
We cast this problem into the AC framework by considering molecular dynamics (MD) simulations and experimental results as separate fidelities (low and high, respectively), with an outer loop problem of optimizing the process parameters for material quality. We then use the AC framework to learn the underlying connection between the simulation and experiment, while taking into account experimental uncertainty with limited samples.
@@
For illustration, at NREL, we are using the Adaptive Mesh and Algorithm Refinement (AMAR) approach. It allows for active coupling between mesh-based partial differential equation (PDE) solvers at the continuum level along with particle-based solvers at the molecular level. The computational domain capturing the physics at the reactor and device levels is discretized using a background mesh. At locations of governing physical phenomena, e.g., catalyst/electrode surfaces, the mesh is refined using adaptive mesh refinement (AMR) and the continuum solver is replaced with a different physical approach (e.g., kinetic Monte-Carlo (KMC) simulations), referred to as algorithm refinement. Information is transferred back and forth between the continuum and particle solvers in real-time. This has been used to model the dynamics inside a chemical vapor deposition reactor for perovskite crystal growth. The continuum flow solver transports species to a substrate. Several parallel KMC simulations are then initiated for each location on the substrate with its corresponding gas phase compositions, pressure and temperature. The net flux of species at the surface is coupled to the continuum as a boundary condition.
@@
The AC framework can also be used for engineering optimization for the design of the macro-scale conditions in the chemical vapor deposition reactor (e.g., ambient pressure, temperature, species composition, jet configuration), which determine the quality of the perovskite crystal grown at the substrate (e.g., number of defects, density, layer compositions, roughness). The AC framework is also useful for managing the uncertainty arising from coupling a deterministic continuum model with a stochastic particle model. Tracking the effect of the resulting uncertainties on the quantity of interest (QOI) (e.g., crystal layer growth, response of the flow at the substrate) is of vital importance.
@@ Acknowledgements
We would like to acknowledge the domain experts who supported the integration of AC with our application areas. Nick Wunder and Monte Lunacek developed the HERO framework and aided with the integration with AC. Davi Febba Marcelo, Stephen Schaefer, and Andriy Zakutayev developed the experimental infrastructure for the material synthesis portion, while Aidan Raver and Garritt Tucker provided simulation support.
@@
Kevin Patrick Griffin is Postdoctoral Research at NREL in Golden, CO, 80401, USA. His research interests include data-driven modeling, turbulence, and computational fluid dynamics (CFD). Dr. Griffin received his Ph.D.  in Mechanical Engineering from Stanford University. Contact him at kevin.griffin@nrel.gov.
@@
Ross Larsen is a Senior Scientist at NREL in Golden, CO, 80401, USA. His research interests include electronic structure and dynamics in condensed phases, statistical mechanics of disordered systems, and molecular dynamics and Monte Carlo simulation methods. Larsen received his Ph.D. in Physics from Brown University. Contact him at ross.larsen@nrel.gov.
@@
Jibonananda Sanyal is a Group Manager for Hybrid Energy Systems at NREL. His research interests include renewable energy, end-uses of energy, and applications of computing. Sanyal received his PhD in Computer Science from Mississippi State University. He is a Senior Member of IEEE. Contact him at jibo.sanyal@nrel.gov.
Hariswaran Sitaraman is a computational scientist at NREL at Golden, CO, 80401, USA. His research interests include development of performance portable solvers and simulations of multiphase reacting and electrical driven flows. Sitaraman received his Ph.D. in Aerospace Engineering, University of Texas at Austin. Contact him at hariswaran.sitaraman@nrel.gov.
@@
Ethan Young is a Computational Scientist at NREL at Golden, CO, 80401, USA. His research interests include developing high-performance CFD tools for fluid-structure interaction problems, porous multiphase flows, and design and control optimization problems. Young received his Ph.D. in Biomedical Engineering from the University of California, Los Angles. Contact him at ethan.young@nrel.gov.